OpenStudy (anonymous):

Which ratio is not equal to 9/12. a.18/24 b.3/4 c.15/20 d.1827

OpenStudy (anonymous):

18/27

OpenStudy (anonymous):

9/12 = 0.75 18/24 = 0.75 3/4 = 0.75 15/20=0.75 18/27 = 0.666 or something :d

OpenStudy (anonymous):

thnx